# Observability conditions for neural state-space models with eigenvalues and their roots of unity

### v1

@@ 3.2 Loss from the Hautus lemma
It is notable to us that this method is equivalent to ensuring $Cv\neq 0$ for all eignevectors $v$ of $A$. This is because $Av-\lambda Iv=0$, therefore the matrix can never be full column rank if $v$ is in the kernel of $C$. Indeed, we can improve efficiency of this by diagonalizing $A$ and then taking the loss
$\mathcal{L}=\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-||Cv_{j}||_{2}^{2}\Bigg{)}.$ (27)
Here, $v_{j}$ are the eigenvectors of $A$. The computational cost of this is $O(mn^{2}+mn^{2}+mn)$, from the $CV$ formulation, the multiplication, and the norm respectively. As we will see, we can improve this by not needing to compute $CV$ but instead only needing light conditions on the eigenvalues, thus our methods are superior.
@@ 4.2 Observability with the Fourier transform
Then with high probability the system is observable.
@@ 4.3 Observability with a more applicable Fourier transform
$+\sum_{j_{1}}\sum_{j_{2}\neq j_{1}}\sum_{k}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\Psi_{kk}^{j_{1}}-\Psi_{kk}^{j_{2}}\Bigg{|}^{2}\Bigg{)}.$ (48)
@@
Figure 3: We illustrate that Theorem 4 is empirically valid. (a) illustrates observability and the Fourier and eigenvalue loss as in equation 205 (b) illustrates observability loss with the observability matrix determinant. We choose $n=50,m=25$. We take the log of the determinant, which helps nondegeneracy from the accumulation of small values. If the observability matrix were low rank, the blue loss would be infinite, which we do not have (except at initialization only), so the system is observable. As we can see, loss achieves exactly zero from 205 because we use a relu-type loss.
Figure 4: In (a), regarding Theorem 4, as we can see, satisfying the distinct eigenvalues condition does not imply the distinctiveness of $\Psi$ with respect to $j$ is automatically satisfied. Thus, both loss terms of equation 205 are necessary to enforce observability with high probability. In (b), we provide a Hautus loss condition and show that the minimum norm on a column of $CV$ is nonzero regarding Theorem 4 in a very simple state-space model setup with learnable initial hidden state. Here, $n=50,m=25$ as well.
@@
$\mathcal{L}=\sum_{j}\sum_{i=2}^{n}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\tilde{C}_{1i}e^{j\Delta\lambda_{i}}-\tilde{C}_{1(i-1)}e^{j\Delta\lambda_{i-1}}\Bigg{|}\Bigg{)},$ (51)
@@
$\mathcal{L}=\sum_{k}\sum_{i=2}^{n}\text{relu}\Bigg{(}\text{positive constant}-\sum_{j}\Bigg{|}\tilde{C}_{ji}e^{k\Delta\lambda_{i}}-\tilde{C}_{j(i-1)}e^{k\Delta\lambda_{i-1}}\Bigg{|}\Bigg{)}.$ (52)
@@ 4.6 Algorithmic convergence and training
Theorem 6. Suppose $f$ decays at a rate
@@ B.1 Observability with the Fourier transform with the impulse response
$\varphi_{j,j_{2}}=\text{column}\Bigg{(}\frac{-[(Q^{j})^{-1}\tilde{P}^{j}]_{1}}{1-e^{\lambda_{1}\Delta t-2\pi i{j_{2}}/L}},\ldots,\frac{-[(Q^{j})^{-1}\tilde{P}^{j}]_{m}}{1-e^{\lambda_{m}\Delta t-2\pi i{j_{2}}/L}},\frac{1}{1-e^{\lambda_{1}\Delta t-2\pi i{j_{2}}/L}},\ldots,\frac{1}{1-e^{\lambda_{n}\Delta t-2\pi i{j_{2}}/L}}\Bigg{)}.$ (99)
@@
Proof. Observe for general $j>0$,
@@
$+\sum_{j_{1}}\sum_{j_{2}\neq j_{1}}\sum_{k}\text{relu}\Bigg{(}\text{positive constant}-\Big{|}\Phi_{kk}^{j_{1}}-\Phi_{kk}^{j_{2}}\Big{|}\Bigg{)}.$ (121)
@@ B.2 Observability using the Fourier transform with convolutional setup
$+\sum_{j_{1}}\sum_{j_{2}\neq j_{1}}\sum_{k}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\Psi_{kk}^{j_{1}}-\Psi_{kk}^{j_{2}}\Bigg{|}^{2}\Bigg{)}.$ (169)
@@
Claim. Let $\Psi^{j}$ be the diagonal matrix defined in 201. Let $C\in\mathbb{R}^{m\times n}$ be full row rank and wide, and let $V\in\mathbb{R}^{n\times n}$ be surjective. Then equation 204 is satisfied with high probability.
Sketch of proof. The argument here is that $j$ is coupled nonlinearly with the eigenvalues of $\lambda^{k}$. This nonlinearity is more rich than simply exponentiating the eigenvalues of the form $\Lambda^{k}$, where $k$ is an exponent, and has greater probability to shift the eigenspace across $j$ via learning. More specifically, eigenvalue exponentiation causes uniform scaling in the complex plane, since $(re^{i\theta})^{k}=r^{k}e^{ik\theta}$ simply scales the radius and multiplies the angle the same across $\theta$. It is the same radial increment per power. On the other hand, our matrix $\Psi^{j}$ has intricate scaling, and the eigenvalues shift and disperse among the complex plane in a non-uniform manner. This helps contribute towards rowspace intersections of dimension less than $m$. It is with greater probability that $\Psi^{j}$ shifts the rowspace rather than exponentiating $\Lambda^{k}$. We demonstrate this phenomenon empirically.
@@
This will in term promote diverseness and help facilitate the permutation-like qualities of $\Psi^{j}$.
@@
Figure 8: We illustrate the same phenomenon as Figure 6 again. Here, the eigenvalues lie among the upper half of the complex plane only. We choose a radius centered at $r=0$ and sample with a unit normal scaled by $0.5$. Again, what we propose is superior.
@@ Appendix C Observability for Mamba: an adapation of the Hautus eigenvector condition
$\mathcal{L}=\sum_{j}\sum_{i=2}^{n}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\tilde{C}_{1i}e^{j\Delta\lambda_{i}}-\tilde{C}_{1(i-1)}e^{j\Delta\lambda_{i-1}}\Bigg{|}\Bigg{)},$ (217)
@@
to be full column rank. This is very simple to enforce as full column rank. We only require that
$\tilde{C}_{1i}e^{j\Delta\lambda_{i}}\neq\tilde{C}_{1(i-1)}e^{j\Delta\lambda_{i-1}}.$ (229)
Thus our loss function is
$\mathcal{L}=\sum_{j}\sum_{i=2}^{n}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\tilde{C}_{1i}e^{j\Delta\lambda_{i}}-\tilde{C}_{1(i-1)}e^{j\Delta\lambda_{i-1}}\Bigg{|}\Bigg{)},$ (230)
@@
$\mathcal{L}=\sum_{k}\sum_{i=2}^{n}\text{relu}\Bigg{(}\text{positive constant}-\sum_{j}\Bigg{|}\tilde{C}_{ji}e^{k\Delta\lambda_{i}}-\tilde{C}_{j(i-1)}e^{k\Delta\lambda_{i-1}}\Bigg{|}\Bigg{)}.$ (231)
@@
However, we remark we found the exponentiation essentially exactly equivalent in computation time if the procedure is used to assign the exponential $z,y\leftarrow e^{x}$, and then $z\leftarrow zy$ is taken. Observe taking $z\leftarrow zy$ is a better strategy for computing the powers too, otherwise all powers $1,\ldots,k$ would need to be taken altogether. Advantages of using exponentials include qualities such as smoother and nonzero gradients and better-conditioning.
@@ Appendix D Codependence: parameter reduction with efficient exponentiation for observability
D.1 Training algorithm
Theorem 6. Suppose $f$ decays at a rate

### v2

@@ 3.2 Loss from the Hautus lemma
It is notable to us that it is necessary $Cv\neq 0$ for all eigenvectors $v$ of $A$. This is because $Av-\lambda Iv=0$, therefore the matrix can never be full column rank if $v$ is in the kernel of $C$. Indeed, we can improve efficiency of this by diagonalizing $A$ and then taking the loss
=$\mathcal{L}\stackrel{{\scriptstyle!}}{{=}}\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-||Cv_{j}||_{2}^{2}\Bigg{)}.$ (27)
Here, $v_{j}$ are the eigenvectors of $A$. We remark this condition is necessary, not sufficient, and so we have denoted that this loss function is incomplete for observability. Consider the counterexample $A=I=I^{3},C=\text{concat}\{(1\ 0\ 1),(0\ 1\ 1)\}$, in which no eigenvector is mapped to triviality but the system is clearly not observable. The computational cost of this is $O(mn^{2}+mn)$, from the $CV$ formulation and the norm respectively. As we will see, we can improve this by not needing to compute $CV$ but instead only needing light conditions on the eigenvalues, thus our methods are superior.
@@ 4.2 Observability with the Fourier transform
$+\text{relu}\Bigg{(}\Bigg{|}\frac{1}{n}\sum_{k}e^{i\theta_{k}(\Lambda)}\Bigg{|}-\text{positive constant}\Bigg{)}$ (42)
@@
$\theta_{k}(\Lambda)$ denotes the angular part of the complex exponentials of the entries of $\Lambda$. Then with high probability the system is observable.
@@ 4.3 Observability with a more applicable Fourier transform
Figure 3: This figure highlights a few important concepts. First, this figure illustrates the number of distinct kernels across $\Psi^{j}$ with respect to $j$, with similarity to the figures in the appendix. Higher is better, so our method clearly outperforms. The primary concept this figure illustrates is that the loss functions we develop in our Fourier-based theorems have significance, and they outperform eigenvalues that are simply random in ensuring the distinct kernel conditions hold. Here, green corresponds to the eigenvalues learned with our custom loss functions, while red corresponds to random eigenvalues. A second concept this figure illustrates is that we can control $\Delta$ predictably. In particular, the same $\Delta$ used in the loss function is the same $\Delta$ used to recreate these kernel results. We chose $\Delta=0.1$, $n=40,m=25.$
@@
$+\sum_{j_{1}}\sum_{j_{2}\neq j_{1}}\sum_{k}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\Psi_{kk}^{j_{1}}-\Psi_{kk}^{j_{2}}\Bigg{|}^{2}\Bigg{)}$ (49)
$+\text{relu}\Bigg{(}\Bigg{|}\frac{1}{n}\sum_{k}e^{i\theta_{k}(\Lambda)}\Bigg{|}-\text{positive constant}\Bigg{)}.$ (50)
@@
Remark. The choice of positive constants in the loss can vary between the two terms, and this will impact performance.
@@
Figure 4: We illustrate that Theorem 4 is empirically valid. (a) illustrates observability and the Fourier and eigenvalue loss as in equation 48 (b) illustrates observability loss with the observability matrix determinant. We choose $n=50,m=25$. We take the log of the determinant, which helps nondegeneracy from the accumulation of small values. If the observability matrix were low rank, the blue loss would be infinite, which we do not have (except near initialization only), so the system is observable. As we can see, loss achieves exactly zero from 48 because we use a relu-type loss.
Figure 5: In (a), regarding Theorem 4, as we can see, satisfying the distinct eigenvalues condition does not imply the distinctiveness of $\Psi$ with respect to $j$ is automatically satisfied. Thus, both of these loss terms of equation 48 are valuable in enforcing observability with high probability. In (b), we provide a Hautus loss condition and show that the minimum norm on a column of $CV$ is nonzero regarding Theorem 4 in a very simple state-space model setup with learnable initial hidden state. Here, $n=50,m=25$ as well. Again, this eigenvector condition is necessary, not sufficient, as we provided a counterexample in section 3.2.
We make some concluding remarks on this theory. It is still necessary $Cv\neq 0$ is valid for the result to hold, which is the Hautus lemma. Instead, the eigenvalues act more of a guide to help the permutation structure of the observability matrix. For example, the observability matrix $(CV^{T}\ CV^{T}\ CV^{T}\ \ldots)^{T}$ still cannot be full column rank because $C$ is wide, where we have diagonalized $A$ and the eigenvalue matrix is the identity. Thus, eigenvalues work at the interplay of these systems, and the high probability necessity holds.
@@
$\mathcal{L}=$ relu(positive $\ \text{relu}\Bigg{(}\text{positive constant}-\min_{k_{1},k_{2}}\Bigg{|}\lambda_{k_{1}}-\lambda_{k_{2}}\Bigg{|}\Bigg{)}+\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-|\tilde{C}_{1j}|\Bigg{)}$ (53)
@@
$\mathcal{L}=$ relu(positive $\ \text{relu}\Bigg{(}\text{positive constant}-\min_{k_{1},k_{2}}\Bigg{|}\lambda_{k_{1}}-\lambda_{k_{2}}\Bigg{|}\Bigg{)}+\sum_{i}\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-|\tilde{C}_{ij}|\Bigg{)}.$ (54)
@@ 4.6 Algorithmic convergence and training
Theorem 6. Suppose loss $f$ decays at a rate
@@ B.1 Observability with the Fourier transform with the impulse response
$+\text{relu}\Bigg{(}\Bigg{|}\frac{1}{n}\sum_{k}e^{i\theta_{k}(\Lambda)}\Bigg{|}-\text{positive constant}\Bigg{)}$ (98)
@@
$\varphi_{j,j_{2}}=\text{column}\Bigg{(}\frac{-[(Q^{j})^{-1}\tilde{P}^{j}]_{1}}{1-e^{\lambda_{1}\Delta t-2\pi i{j_{2}}/L}},\ldots,\frac{-[(Q^{j})^{-1}\tilde{P}^{j}]_{m}}{1-e^{\lambda_{m}\Delta t-2\pi i{j_{2}}/L}},\frac{1}{1-e^{\lambda_{1}\Delta t-2\pi i{j_{2}}/L}},\ldots,\frac{1}{1-e^{\lambda_{n}\Delta t-2\pi i{j_{2}}/L}}\Bigg{)}$ (102)
@@
Proof. First, we remark this proof goes hand-in-hand with Theorem 4, so we also advise reading the proof of that theorem if details seem omitted.
Observe for general $j>0$,
@@
$+\text{relu}\Bigg{(}\Bigg{|}\frac{1}{n}\sum_{k}e^{i\theta_{k}(\Lambda)}\Bigg{|}-\text{positive constant}\Bigg{)}.$ (125)
@@
The last term involving variance ensures dispersion of $\Phi$ across the complex plane.
@@ B.2 Observability using the Fourier transform with convolutional setup
$+\sum_{j_{1}}\sum_{j_{2}\neq j_{1}}\sum_{k}\text{relu}\Bigg{(}\text{positive constant}-\Bigg{|}\Psi_{kk}^{j_{1}}-\Psi_{kk}^{j_{2}}\Bigg{|}^{2}\Bigg{)}$ (173)
$+\text{relu}\Bigg{(}\Bigg{|}\frac{1}{n}\sum_{k}e^{i\theta_{k}(\Lambda)}\Bigg{|}-\text{positive constant}\Bigg{)}.$ (174)
@@
Claim. Let $\Psi^{j}$ be the diagonal matrix defined in 206. Let $C\in\mathbb{R}^{m\times n}$ be full row rank and wide, and let $V\in\mathbb{R}^{n\times n}$ be surjective. Then equation 209 is satisfied with high probability.
Sketch of proof. The argument here is that $j$ is coupled nonlinearly with the eigenvalues of $\lambda^{k}$. This nonlinearity is more rich than simply exponentiating the eigenvalues of the form $\Lambda^{k}$, where $k$ is an exponent, and has greater probability to shift the eigenspace across $j$ via learning. More specifically, eigenvalue exponentiation causes uniform scaling in the complex plane, since $(re^{i\theta})^{k}=r^{k}e^{ik\theta}$ simply scales the radius and multiplies the angle the same across $\theta$. It is the same radial increment per power. On the other hand, our matrix $\Psi^{j}$ has intricate scaling, and the eigenvalues shift and disperse among the complex plane in a non-uniform manner. This helps contribute towards rowspace intersections of dimension less than $m$. It is with greater probability that $\Psi^{j}$ shifts the rowspace rather than exponentiating $\Lambda^{k}$. Moreover, we specifically constrain the eigenvalues to enforce this property. We have found the losses of
  • (Sufficiently up to error tolerances) (D)istinct eigenvalues
  • Sufficient distinctness across $\Psi^{j}$ with respect to $j$
  • Sufficient angular diversity across the eigenvalues
work well. Intuitively, the first two are not hard to understand. The third one is the hard one, and indeed, we discovered this through use of experimentation (see Figure 3). We demonstrate the phenomenon that $\Psi^{j}$ promotes kernel diversity empirically.
@@
This will in term promote diverseness and help facilitate the permutation-like qualities of $\Psi^{j}$. Moreover, to ensure the dispersion of eigenvalues across the complex plane, we can also penalize radial differences in the complex exponents of $re^{i\theta}$. We also develop the loss
$\mathcal{L}=\text{relu}\Bigg{(}\Bigg{|}\frac{1}{n}\sum_{k}e^{i\theta_{k}(\Lambda)}\Bigg{|}-\text{positive constant}\Bigg{)}.$ (211)
Here, we denote $\theta(\cdot)$ the angular part of the complex number as an exponential. This loss function works by omitting the radial aspect of the eigenvalues, thus by only including the angles with a sense of normalization, angular diversity should lead to a balancing among the vectors. We desire the absolute summation close to zero. We have found empirical success in choosing radial diversity among the eigenvalues rather than across $j$ with $\Psi^{j}$.
@@ Appendix C Observability for Mamba: an adapation of the Hautus eigenvector condition
$\mathcal{L}=$ relu(positive $\ \text{relu}\Bigg{(}\text{positive constant}-\min_{k_{1},k_{2}}\Bigg{|}\lambda_{k_{1}}-\lambda_{k_{2}}\Bigg{|}\Bigg{)}+\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-|\tilde{C}_{1j}|\Bigg{)}$ (223)
@@
to be full column rank. This is very simple to enforce as full column rank. We only need the eigenvalues are distinct and all $\tilde{C}_{1j}$ are nonzero. Thus our loss function is
$\mathcal{L}=$ relu(positive $\ \text{relu}\Bigg{(}\text{positive constant}-\min_{k_{1},k_{2}}\Bigg{|}\lambda_{k_{1}}-\lambda_{k_{2}}\Bigg{|}\Bigg{)}+\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-|\tilde{C}_{1j}|\Bigg{)},$ (235)
@@
$\mathcal{L}=$ relu(positive $\ \text{relu}\Bigg{(}\text{positive constant}-\min_{k_{1},k_{2}}\Bigg{|}\lambda_{k_{1}}-\lambda_{k_{2}}\Bigg{|}\Bigg{)}+\sum_{i}\sum_{j}\text{relu}\Bigg{(}\text{positive constant}-|\tilde{C}_{ij}|\Bigg{)}.$ (236)
@@
This strategy does not change the loss function.
@@ Appendix D Codependence: parameter reduction with efficient exponentiation for observability
Figure 13: We illustrate results upon our proposed algorithm versus a more traditional algorithm. In this experiment, we use a Mamba architecture on basic sinusoidal sequences, and replace parameters with those we create. In particular, our algorithm yields more stable (stable in the sense of smaller in value, not in the sense of stochasticity) gradients with respect to $B,\mathcal{B}$, whereas a vanilla algorithm yields a gradient that is much higher. We mention training loss across both methods is relatively consistent, thus we remark the diminished gradient of our method is both stable and does not interfere with the training process. Our method is the gradient $\alpha\nabla_{\mathcal{B}}\mathcal{L}$ at a new step where $\alpha=||\mathcal{B}_{k}-\mathcal{B}_{k-1}||_{F}^{1/q}$ is the appropriate scaling for choice of $k$, and gradients at older steps are frozen. Gradients for the classical algorithm are the per step gradient $\nabla_{\mathcal{B}}\mathcal{L}$. In particular, our method slowly smooths into the correct parameter, thus is predictable and stable, whereas the gradients maintain large value in a more baseline method.
Figure 14: We illustrate training loss with our proposed algorithm versus a classical objective.
Theorem 6. Suppose loss $f$ decays at a rate